If 5.50 kJ if he is applied to a 0.5 85 MOL sample of solid copper at 25 Celsius what will the coppers final temperature be in Celsius the specific heat of solid copper is 0.385 J/G•K


1
Expert's answer
2021-03-11T08:41:07-0500

"Q=mc\theta"


"1mol=65g"

"0.585moles=?"


"(\frac{65×0.585}{1})g=25.025g"


5.50=25.025g×0.385×(x-25)

5.50=9.634625x"-" 246.865625


"\frac{246.8656}{9.634625}=\frac{9.64625x}{9.634625}"


"x=25.57\u00b0C"